# Cost-Sensitive Learning to Defer to Multiple Experts with Workload Constraints

### v2

@@ 4.1 Experimental Setup
As the base dataset, we choose to use the publicly available bank-account-fraud dataset [Jesus et al., 2022] (Version 1). This tabular dataset is comprised of one million synthetically generated bank account opening applications, where the label denotes whether the instance is a fraudulent (1) or a legitimate (0) application. The features of each instance contain information about the application and the applicant, and the task of a decision maker (automated or human) is to either accept (0) or reject (1) it.
These applications were generated based on anonymized real-world bank account applications, and, as such, this dataset poses challenges typical of real-world high-stakes applications. Firstly, there is a high class imbalance, with 1.1% fraud prevalence over the entire dataset. Furthermore, there are changes in the data distribution over time, often referred to as concept drift [Gama et al., 2014], which can severely impact the predictive performance of ML Models.
@@
This is a cost-sensitive task, where the cost of a false positive (incorrectly rejecting a legitimate application) must be weighed against the cost of a false negative (incorrectly accepting a fraudulent application). Due to the low fraud prevalence, metrics such as accuracy are not adequate to measure the performance of both ML models and deferral systems. The optimization objective used by Jesus et al. [2022] is a Neyman-Pearson criterion, in this case, maximizing recall at 5% false positive rate (FPR), which establishes an implicit relationship between the costs of false positive and a false negative errors. However, for the cost-sensitive learning method described in Section 3.4, we need to have access to the cost structure of the task at hand.
@@
In a cost-sensitive task, the optimization goal is to obtain a set of predictions $\hat{y}$ that minimize the expected misclassification cost $\mathbb{E}[\mathcal{C}]$. Assuming that correct classifications carry no cost, the relevant parameter is the ratio $\lambda=c_{FP}/c_{FN}$, were $c_{FP}$ and $c_{FN}$ are the costs of false positive and false negative errors, respectively. The objective is thus
$\frac{1}{N}\sum_{i=1}^{N}\big{[}\lambda\mathbb{I}[\mbox{y}_{i}=0\land\hat{y}_{i}=1]+\mathbb{I}[\mbox{y}_{i}=1\land\hat{y}_{i}=0]\big{]}.$ (11)
Minimizing this quantity is equivalent to minimizing the average cost, as division by a constant will not affect the ranking of different assignments.
@@
As such, all that remains to establish is a relationship between the Neyman-Pearson criterion and the value of $\lambda$. To to do so, we follow the approach detailed in Section A.2 of the Appendix, which yields the theoretical value $\lambda_{t}=0.057$. To test performance under variable cost structures, we will conduct experiments for the values $\lambda\in\{\lambda_{t}/5,\lambda_{t},5\lambda_{t}\}$. These alternative scenarios are not strictly comparable, as if the cost structure were $\lambda\in\{\lambda_{t}/5,5\lambda_{t}\}$, the optimization of the Alert Model would not correspond to maximizing the recall at 5% FPR. Nevertheless, we choose to introduce these variations to test the impact on system performance of changing the cost ratio $\lambda$.
Alert Review Setup
There are several possible ways for models and humans to cooperate. In previous L2D research, it is a common assumption that any instance can be deferred to either the model or the expert team. However, in real-world settings, due to limitations in human work capacity, it is common to use an ML model to screen instances, raising alerts that are then reviewed by human experts [De-Arteaga et al., 2020, Han et al., 2020]. In an alert review setting, humans only predict on a fraction of the feature space, that is, the instances flagged by the Alert Model. We will train a L2D system to work in tandem with the Alert Model, by deferring the alerts in an intelligent manner. We calculate the Alert Model’s alert rate $a_{r}$, that is, the fraction of instances flagged for human review, by determining the FPR of the Alert Model in validation. We create 6 distinct alert review scenarios by varying the alert rate $a_{r}\in\{5\%FPR,15\%FPR\}$ for each of the 3 values of $\lambda$ detailed in the previous paragraph.
Alert Model Training
We train the Alert Model to predict the fraud label on the first three months of the dataset, validating its performance on the fourth month. We use the LightGBM [Ke et al., 2017] algorithm, due to its proven high performance on tabular data [Shwartz-Ziv and Armon, 2022, Borisov et al., 2022]. Details on the training process of the classifier are given in Section A.1 of the Appendix.
@@
Our expert generation approach is based on instance-dependent noise, in order to obtain more realistic experts, whose probability of error varies with the properties of each instance. We generate synthetic predictions by flipping each label $\mbox{y}_{i}$ with probability $\mathbb{P}(m_{j,i}\neq\mbox{y}_{i}\lvert\bm{x}_{i},\mbox{y}_{i})$. In some HAIC systems, the model score for a given instance may also be shown to the expert [Amarasinghe et al., 2022, De-Arteaga et al., 2020, Levy et al., 2021], so an expert’s decision may also be dependent on an ML model score $m(\bm{x}_{i})$. We define the expert’s probabilities of error, for a given instance, as a function of a pre-processed version of its features $\bm{\bar{x}}_{i}$ and the Alert Model’s score $M(\bm{x}_{i})$, so that the feature scale does not impact the relative importance of each quantity. The probabilities of error are given by
@@
$\begin{gathered}\begin{cases}\mathbb{P}(m_{j,i}=1\lvert\mbox{y}_{i}=0,\bm{x}_{i})=\sigma\Big{(}\beta_{0}-\alpha\frac{\bm{w}\cdot\bm{\bar{x}}_{i}+w_{M}M(\bm{x}_{i})}{\sqrt{\bm{w}\cdot\bm{w}+w_{M}^{2}}}\Big{)}\\ \mathbb{P}(m_{j,i}=0\lvert\mbox{y}_{i}=1,\bm{x}_{i})=\sigma\Big{(}\beta_{1}+\alpha\frac{\bm{w}\cdot\bm{\bar{x}}_{i}+\bm{w}_{M}M(\bm{x}_{i})}{\sqrt{\bm{w}\cdot\bm{w}+w_{M}^{2}}}\Big{)},\end{cases}\end{gathered}$ (12)
@@
where $\sigma$ denotes a sigmoid function. Each expert’s probabilities of the two types of error are parameterized by five parameters: $\beta_{0},\beta_{1},\alpha,\bm{w}$ and $w_{M}$. The weight vector $\bm{w}$ embodies a relation between the features and the probability of error. The feature weights are normalized so that we can separately control, via $\alpha$, the overall magnitude of the dependence of the probability of error on the instance’s features. The values of $\beta_{1}$ and $\beta_{0}$ control the base probability of error. The motivation for this approach is explained further in section C.1 of the Appendix.
We can calculate the expected cost resulting from an expert’s decisions as
@@
$\begin{split}\mathbb{E}[\mathcal{C}]_{j}=\mathbb{E}_{\mbox{\scriptsize{}y}}&\big{[}\lambda\mathbb{P}(m_{j}=1\land\mbox{y}=0)+\mathbb{P}(m_{j}=0\land\mbox{y}=1)\big{]}\\ \approx\frac{1}{N}\sum_{i=1}^{N}&\big{[}\lambda\mathbb{P}(m_{j,i}=1\lvert\mbox{y}_{i}=0)\mathbb{P}(\mbox{y}_{i}=0)+\mathbb{P}(m_{j,i}=0\lvert\mbox{y}_{i}=1)\mathbb{P}(\mbox{y}_{i}=1)\big{]}.\end{split}$ (13)
@@
For our setup to be realistic, we assume an expert’s decisions must, on average, incur a lower cost than simply automatically rejecting all flagged transactions. Otherwise, assuming random assignment, having that expert in the human team would harm the performance of the system as a whole. As the expert’s average misclassification cost is dependent on the prevalence $\mathbb{P}(\mbox{y}=1)$ and cost structure as defined by $\lambda$, a team of experts with the exact same parameters will have different expected misclassification costs, depending on the alert review scenario in question. For this reason, in each of the six aforementioned settings, we generate a different team of 9 synthetic experts by first fixing their feature weights $\bm{w}$, and sampling their expected misclassification cost. Then, we calculate the values of $\beta_{0},\beta_{1}$ that achieve the sampled misclassification cost, thus obtaining a team of complex synthetic experts with desirable properties within each scenario. Further details on the sampling and expert generation process, as well as a description of each of the expert teams’ properties are available in Sections C.2 and C.3 of the Appendix.
@@
To generate realistic expert data availability, we assume that the Alert Model is deployed in months 4-7, and that each alert is deferred to a randomly chosen expert. This generates a history of expert decisions with only one expert’s prediction per instance. To introduce variability to the L2D training process, the random distribution of cases throughout experts was done with 5 different random seeds per scenario. For each of the 3 $5\%$ FPR alert rate settings, this results in a set of 2.9K predictions per expert. So that all settings have the same amount of data, for the $15\%$ FPR alert rate scenarios we also sample 2.9K predictions from each of the 10 experts.
@@
In our experiments, we want to reliably measure the ability of our method to optimize the distribution of instances throughout the human team and the classifier $h$. In real-world scenarios, the cost of querying $h$ is much lower than that of querying a human expert, as the work capacity of the classifier is virtually limitless. However, in order to ensure that the assignment systems are able to reliably model the human behavior of all experts, we will force the assigner to distribute the instances throughout the expert team and the classifier $h$ in equal amounts. As the expert teams are comprised of 9 experts, this means that a tenth of the test set will be deferred to each expert/classifier. To introduce variability, we also create four distinct test settings where each expert/classifier has a different capacity, by sampling the values from $\mathcal{N}(N_{test}/10,N_{test}/50)$. As such, for each scenario, defined by the $(a_{r},\lambda)$ pair, there are a total of 25 test variations, which result from combining each of the 5 training seeds with the 5 different capacity constraint settings.
@@ 4.2 Baselines
For an L2D baseline, we use the multi-expert learning to defer OvA algorithm, proposed by Verma et al. [2023]. This method originally takes training samples of the form $D_{i}=\{\boldsymbol{x}_{i},\mbox{y}_{i},m_{i,1},...,m_{j,i}\}$ and assumes the existence of a set of every expert’s predictions for each training instance; however, this is not strictly necessary.
@@
$\Psi_{\mbox{\scriptsize OvA}}=$ $\Phi[g_{y}(\boldsymbol{x})]+\sum_{y^{\prime}\in\mathcal{Y},y^{\prime}\neq y}\Phi[-g_{y}^{\prime}(\boldsymbol{x})]+\sum_{j=1}^{J}\Phi[-g_{\perp,j}(\boldsymbol{x})]+$
$\sum_{j=1}^{J}\mathbb{I}[m_{j}=y](\Phi[g_{\perp,j}(\boldsymbol{x})]-\Phi[-g_{\perp,j}(\boldsymbol{x})]),$
where $\Phi:\{\pm 1\}\times\mathbb{R}\rightarrow\mathbb{R}_{+}$ is a strictly proper binary surrogate loss. (The authors also propose using the logistic loss.) Verma et al. [2023] then prove that the minimizer of the pointwise inner risk of this surrogate loss can be analyzed in terms of the pointwise minimizer of the risk for each of the $K+J$ underlying OvA binary classification problems, concluding that the minimizer of the pointwise inner $\Psi_{\mbox{\scriptsize OvA}}$-risk, $\boldsymbol{g}^{*}$, is comprised of the minimizer of the inner $\Phi$-risk for each $i$th binary classification problem, $g_{i}^{*}$. As such, in a scenario where only one expert’s prediction is associated with each instance, each binary classifier $g_{\perp,j}$ can be trained independently of the others. By training each binary classifier $g_{\perp,j}$ with the subset of the training sample containing expert $j$’s predictions, we obtain the best possible estimates of each pointwise inner $\Phi$-risk minimizer $g_{i}^{*}$ given the available data. To adapt the OvA method to a cost sensitive scenario, we can again use the rescaling approach detailed in section 3.4, minimizing the expected misclassification cost.
The OvA method does not support assignment with capacity constraints, as such, we proceed similarly to Verma et al. [2023], by considering the maximum out of the rejection classifiers’ predictions. Should the capacity of the selected expert be exhausted, deferral is done to the second highest scoring expert, and so on.
@@ 5.1 Classifier h - Performance and Calibration
We first evaluate the predictive performance and calibration of the classifier $h$. We assess how well the classifier is able to rank probabilities of correctness, and then evaluate its calibration, by using the expected calibration error [Guo et al., 2017]. Note that each classifier had a distinct training process, with varying sample weights $c_{i}$ dependent on $\lambda$. As such, these measures must be calculated under the re-weighted data distribution $\tilde{D}$ (see section 3.4). In Table 2, we show that the classifier $h$ is able to rank instances according to their probability of belonging to the positive class.
Table 2: Expected Calibration Error (ECE) and ROC-AUC for Classifier h for all testing scenarios, denoted by the {alert rate ($a_{r}$), cost-structure ($\lambda$)} pairs - Classifier $h$ is able to reliably estimate $\mathbb{P}(\mbox{y}_{i}=1)$

### v3

@@ 4.1 Experimental Setup
As the base dataset, we choose to use the publicly available bank-account-fraud dataset (Jesus et al., 2022) (Version 1). This tabular dataset is comprised of one million synthetically generated bank account opening applications, where the label denotes whether the instance is fraudulent (1) or legitimate (0). The features of each instance contain information about the application and the applicant, and the task of a decision maker (automated or human) is to either accept (0) or reject (1) it.
These applications were generated based on anonymized real-world bank account applications, and, as such, this dataset poses challenges typical of real-world high-stakes applications. Firstly, there is a high class imbalance, with 1.1% fraud prevalence over the entire dataset. Furthermore, there are changes in the data distribution over time, often referred to as concept drift (Gama et al., 2014), which can severely impact the predictive performance of ML Models.
Alert Review Setup
There are several possible ways for models and humans to cooperate. In previous L2D research, it is a common assumption that any instance can be deferred to either the model or the expert team. However, in real-world settings, due to limitations in human work-capacity, it is common to use an Alert Model to screen instances, raising alerts that are then reviewed by human experts (De-Arteaga et al., 2020; Han et al., 2020). In an alert review setting, humans only predict on a fraction of the feature space, that is, the instances flagged by the Alert Model. We will train a L2D system to work in tandem with the Alert Model, by deferring the alerts in an intelligent manner. We calculate the Alert Model’s alert rate $a_{r}$, that is, the fraction of instances flagged for human review, by determining the FPR of the Alert Model in validation. We create distinct alert review scenarios by varying the alert rate $a_{r}\in\{5\%\text{FPR},15\%\text{FPR}\}$.
Alert Model Training
We train the Alert Model to predict the fraud label on the first three months of the dataset, validating its performance on the fourth month. We use the LightGBM (Ke et al., 2017) algorithm, due to its proven high performance on tabular data (Shwartz-Ziv & Armon, 2022; Borisov et al., 2022). Details on the training process of the classifier are given in Section A.1 of the Appendix.
@@
This is a cost-sensitive task, where the cost of a false positive (incorrectly rejecting a legitimate application) must be weighed against the cost of a false negative (incorrectly accepting a fraudulent application). Due to the low fraud prevalence, metrics such as accuracy are not adequate to measure the performance of both ML models and deferral systems. The optimization objective used by Jesus et al. (2022) is a Neyman-Pearson criterion, in this case, maximizing recall at 5% false positive rate (FPR), which establishes an implicit relationship between the costs of false positive and a false negative errors. However, for the cost-sensitive learning method described in Section 3.4, we need to have access to the explicit cost structure of the task at hand.
@@
In a cost-sensitive task, the optimization goal is to obtain a set of predictions $\hat{y}$ that minimize the expected misclassification cost $\mathbb{E}[\mathcal{C}]$. Assuming that correct classifications carry no cost, the relevant parameter is the ratio $\lambda=c_{\mbox{\scriptsize FP}}/c_{\mbox{\scriptsize FN}}$, were $c_{\mbox{\scriptsize FP}}$ and $c_{\mbox{\scriptsize FN}}$ are the costs of false positive and false negative errors, respectively. The objective is thus
$\frac{1}{N}\sum_{i=1}^{N}\big{[}\lambda\mathbb{I}[y_{i}=0\land\hat{y}_{i}=1]+\mathbb{I}[y_{i}=1\land\hat{y}_{i}=0]\big{]}.$ (16)
@@
Minimizing this quantity is equivalent to minimizing the average cost, as division by a constant will not affect the ranking of different assignments. As such, all that remains to be established is a relationship between the Neyman-Pearson criterion and the value of $\lambda$. To to do so, we follow the approach detailed in Section A.2 of the Appendix, which yields the theoretical value $\lambda_{t}=0.057$. To test performance under variable cost structures, we will conduct experiments for the values $\lambda\in\{\lambda_{t}/5,\lambda_{t},5\lambda_{t}\}$. These alternative scenarios are not strictly comparable, as if the cost structure were $\lambda\in\{\lambda_{t}/5,5\lambda_{t}\}$, the optimization of the Alert Model would not correspond to maximizing the recall at 5% FPR. Nevertheless, we choose to introduce these variations to test the impact on system performance of changing the cost ratio $\lambda$. Combining the different alert rates with these values of $\lambda$, we obtain 6 distinct text scenarios.
@@
Our expert generation approach is based on instance-dependent noise, in order to obtain more realistic experts, whose probability of error varies with the properties of each instance. We generate synthetic predictions by flipping each label $y_{i}$ with probability $\mathbb{P}(m_{j,i}\neq y_{i}\lvert\bm{x}_{i},y_{i})$. In some HAIC systems, the model score for a given instance may also be shown to the expert (Amarasinghe et al., 2022; De-Arteaga et al., 2020; Levy et al., 2021), so an expert’s decision may also depend on an ML model score $m(\bm{x}_{i})$. We define the expert’s probabilities of error, for a given instance, as a function of a pre-processed version of its features $\bm{\bar{x}}_{i}$ and the Alert Model’s score $M(\bm{x}_{i})$, so that the feature scale does not impact the relative importance of each quantity. The probabilities of error are given by
@@
$\begin{split}&\mathbb{P}(m_{j,i}=1\lvert y_{i}=0,\bm{x}_{i})=\sigma\Big{(}\beta_{0}-\alpha\frac{\bm{w}\cdot\bm{\bar{x}}_{i}+w_{M}M(\bm{x}_{i})}{\sqrt{\bm{w}\cdot\bm{w}+w_{M}^{2}}}\Big{)}\\ &\mathbb{P}(m_{j,i}=0\lvert y_{i}=1,\bm{x}_{i})=\sigma\Big{(}\beta_{1}+\alpha\frac{\bm{w}\cdot\bm{\bar{x}}_{i}+\bm{w}_{M}M(\bm{x}_{i})}{\sqrt{\bm{w}\cdot\bm{w}+w_{M}^{2}}}\Big{)},\end{split}$ (17)
@@
where $\sigma$ denotes a sigmoid function. Each expert’s probabilities of the two types of error depend on five parameters: $\beta_{0},\beta_{1},\alpha,\bm{w}$, and $w_{M}$. The weight vector $\bm{w}$ embodies a relation between the features and the probability of error. The feature weights are normalized so that we can separately control, via $\alpha$, the overall magnitude of the dependence of the probability of error on the instance’s features. The values of $\beta_{1}$ and $\beta_{0}$ control the base probability of error. The motivation for this approach is explained further in section C.1 of the Appendix.
The expected cost resulting from an expert’s decisions is given by
@@
$\begin{split}\mathbb{E}[\mathcal{C}]_{j}&=\mathbb{E}_{\mbox{\scriptsize{}y}}\big{[}\lambda\mathbb{P}(m_{j}=1\land y=0)+\mathbb{P}(m_{j}=0\land y=1)\big{]}\\ &\approx\frac{1}{N}\sum_{i=1}^{N}\big{[}\lambda\mathbb{P}(m_{j,i}=1\lvert y_{i}=0)\mathbb{P}(y_{i}=0)+\mathbb{P}(m_{j,i}=0\lvert y_{i}=1)\mathbb{P}(y_{i}=1)\big{]}.\end{split}$ (18)
@@
For our setup to be realistic, we assume an expert’s decisions must, on average, incur a lower cost than simply automatically rejecting all flagged transactions. Otherwise, assuming random assignment, having that expert in the human team would harm the performance of the system as a whole. As the expert’s average misclassification cost is dependent on the prevalence $\mathbb{P}(y=1)$ and cost structure as defined by $\lambda$, a team of experts with the exact same parameters will have different expected misclassification costs, depending on the alert review scenario in question. For this reason, in each of the six aforementioned settings, we generate a different team of 9 synthetic experts by first fixing their feature weights $\bm{w}$, and sampling their expected misclassification cost. Then, we calculate the values of $\beta_{0},\beta_{1}$ that achieve the sampled misclassification cost, thus obtaining a team of complex synthetic experts with desirable properties within each scenario. Further details on the sampling and expert generation process, as well as a description of each of the expert teams’ properties are available in Sections C.2 and C.3 of the Appendix.
@@
To generate realistic expert data availability, we assume that the Alert Model is deployed in months 4-7, and that each alert is deferred to a randomly chosen expert. This generates a history of expert decisions with only one expert’s prediction per instance. To introduce variability to the L2D training process, the random distribution of cases throughout experts was done with 5 different random seeds per scenario. For each of the three $5\%$ FPR alert rate settings, this results in a set of 2.9K predictions per expert. So that all settings have the same amount of data, for the $15\%$ FPR alert rate scenarios, we also sample 2.9K predictions from each of the 10 experts.
@@
In our experiments, we want to reliably measure the ability of our method to optimize the distribution of instances throughout the human team and the classifier $h$. In real-world scenarios, the cost of querying $h$ is much lower than that of querying a human expert, as the work-capacity of the classifier is virtually limitless. However, in order to ensure that the assignment systems are able to reliably model the human behavior of all experts, we will force the assigner to distribute the instances throughout the expert team and the classifier $h$ in equal amounts. As the expert teams are comprised of 9 experts, this means that 1/10 of the test set will be deferred to each expert/classifier. To introduce variability, we also create four distinct test settings where each expert/classifier has a different capacity, by sampling the values from $\mathcal{N}(N_{test}/10,N_{test}/50)$. As such, for each scenario defined by the $(a_{r},\lambda)$ pair, there are a total of 25 test variations, which result from combining each of the 5 training seeds with the 5 different capacity constraint settings.
@@ 4.2 Baselines
For an L2D baseline, we use the multi-expert learning to defer OvA algorithm, proposed by Verma et al. (2023). This method originally takes training samples of the form $D_{i}=\{\boldsymbol{x}_{i},y_{i},m_{i,1},...,m_{j,i}\}$ and assumes the existence of a set of every expert’s predictions for each training instance; however, this is not strictly necessary.
@@
$\Psi_{\mbox{\scriptsize OvA}}$ $=\Phi[g_{y}(\boldsymbol{x})]+\sum_{y^{\prime}\in\mathcal{Y},y^{\prime}\neq y}\Phi[g_{y}^{\prime}(\boldsymbol{x})]+\sum_{j=1}^{J}\Phi[-g_{\perp,j}(\boldsymbol{x})]+\sum_{j=1}^{J}\mathbb{I}[m_{j}=y](\Phi[g_{\perp,j}(\boldsymbol{x})]-\Phi[-g_{\perp,j}(\boldsymbol{x})]),$
where $\Phi:\{\pm 1\}\times\mathbb{R}\rightarrow\mathbb{R}_{+}$ is a strictly proper binary surrogate loss. (The authors also propose using the logistic loss.) Verma et al. (2023) then prove that the minimizer of the pointwise inner risk of this surrogate loss can be analyzed in terms of the pointwise minimizer of the risk for each of the $K+J$ underlying OvA binary classification problems, concluding that the minimizer of the pointwise inner $\Psi_{\mbox{\scriptsize OvA}}$-risk, $\boldsymbol{g}^{*}$, is comprised of the minimizer of the inner $\Phi$-risk for each $i$th binary classification problem, $g_{i}^{*}$. As such, in a scenario where only one expert’s prediction is associated with each instance, each binary classifier $g_{\perp,j}$ can be trained independently of the others. By training each binary classifier $g_{\perp,j}$ with the subset of the training sample containing expert $j$’s predictions, we obtain the best possible estimates of each pointwise inner $\Phi$-risk minimizer $g_{i}^{*}$ given the available data. To adapt the OvA method to a cost-sensitive scenario, we can again use the rescaling approach detailed in Section 3.4, minimizing the expected misclassification cost.
The OvA method does not support assignment with capacity constraints, as such, we proceed similarly to Verma et al. (2023), by considering the maximum out of the rejection classifiers’ predictions. If the capacity of the selected expert is exhausted, deferral is done to the second highest scoring expert, and so on.
@@ 5.1 Classifier h - Performance and Calibration
We first evaluate the predictive performance and calibration of the classifier $h$. We assess how well the classifier is able to rank probabilities of correctness, and then evaluate its calibration, by using the expected calibration error (Guo et al., 2017). Note that each classifier had a distinct training process, with varying sample weights $c_{i}$ dependent on $\lambda$. As such, these measures must be calculated under the re-weighted data distribution $\tilde{D}$ (see section 3.4). In Table 2, we show that the classifier $h$ is able to rank instances according to their probability of belonging to the positive class.
Table 2: Expected calibration error (ECE) and ROC-AUC for classifier $h$ for all testing scenarios, denoted by the {alert rate ($a_{r}$), cost-structure ($\lambda$)} pairs – classifier $h$ is able to reliably estimate $\mathbb{P}(y_{i}=1)$